# Explaining word embeddings with perfect fidelity: Case study in research impact prediction

### v1

@@ Results
The word “Remdesivir” on its own has with $\textit{SCORE}(w)=0.961$ also a high value of the predicted significance, although not as high as for imdevimab ($\textit{SCORE}(w)=0.997$), which was thus chosen as the representative for the drug topic.
@@
Table 6: Top words for selected topics. Sorted by the $\textit{SCORE}(w)$, Sim is the cosine similarity of the word vector to the word vector of the topic in the caption of each table, year indicates the earliest publication year in the corpus when the word appears.
Experiment 5: Qualitative evaluation of top-predicted articles
This experiment is based on the same data and predictive model as the previous experiment. However, unlike Experiment 4, which focuses on predictions of entities, this one focuses on the prediction of impactful articles. The predictions are validated through the bibliometric profile of the publishing journals and authors. We analyzed “young” articles published from the year 2021 which have the greatest potential (Score) according to the model.
Figure 4 analyzes a subset of articles with the highest predicted impact (Score), which were published in journals with impact factor. Among the most recent articles (published in 2022), the article that appeared in a journal assigned the highest impact factor value (39.2) is also predicted to have the highest impact according to our method (SMER score). Out of the eight articles with an impact factor, seven appeared in journals assigned an impact factor higher than the median of 4.74 in the source corpus. Note that the impact factor or other journal metadata were not part of the training corpus. Similarly, the authors of most top-ranked articles included an author with an H-index exceeding the median for full professors on journal editorial boards. The high impact factors of journals in which the predicted impactful articles were published, as well as the track-record of their authors, confirm the ability of our model to identify future research impact. The document object identifiers (DOIs) of the specific top-predicted articles are available in the supplementary material.
Figure 4: Top articles by predicted impact (Score) published in 2021. The reference impact factor (blue line) represents the median among the top 10% journals in the category “Health sciences” category. The reference H-index (red line) corresponds to the median H-index of full professors. The attention score and tweets were retrieved from Altmetrics.com.
Discussion
In this section, we present a more detailed examination and discussion of the predicted impactful entities from Experiment 4.
An important property of embeddings is that they can provide predictions even for words that were not contained in the labelled training data for the classification model. This is not possible with models based on the bag of words representation. An example is “nirmatrevil”, which has a high predicted score (0.99) according to our proposed embeddings-based logistic regression model. Since this word only appeared for the first time in an article published in the year 2022, it was not included in the abstracts of any of the articles on which the logistic regression model was trained.
In Table 6(c), we show results for the most important drug-related entities which are similar to the COVID-19 drug imdevimab. To summarize these findings, we note that of the top hits to this drug, seven of these top words are other COVID-19 antibody drugs with the antibody-drug designated suffix -mab. Potentially other methods designed for learning word subtokens could identify further impactful patterns based on word prefix and suffix patterns. However, we observe that the identified important features similar to imdevimab are perfectly categorized as antibody drugs, combinations, or more abstract related terms (’monoclonal’), and include many synonyms. Thus included are: regen-cov (representing the combination of casirivimab (#2) and imdevimab (#1)), ly-cov555 (synonym for bamlanivimab (#4)), regn-cov2 (synonym for casirivimab-imdevimab), ronapreve (synonym for casirivimab-imdevimab), ly-cov016 (also known as etesevimab (#3)), and regn10987 (synonym for imdevimab). In total, including identified synonyms, 8 out of the top important and similar words to imdevimab are synonyms of this drug or its combinations (e.g. the combinations casirivimab-imdevimab, casirivimab-imdevimab, regn-cov2 and ronapreve). The remaining words among the top hits are other COVID-19 antibody drugs, with monoclonal representing the general concept of monoclonal antibodies. With the exception of this general term, all other important words first appeared in the literature in 2020 or later, further validating their relevance for similarity to imdevimab, which is a new drug developed for COVID-19.
In Table 6(d), we also considered similar and important features related to the drug remdevisir, which has been used as a diagnostic word to assess word similarities for COVID-19 literature. In this case, all of the top important features were small-molecule COVID-19 drugs, with one exception being the antibody drug arilumab approved for rheumatoid arthritis and repurposed for COVID-19. We note again that the suffix -ir, indicating an antiviral compound, is found for 8 of the 14 most important words similar to remdesivir, suggesting that subtoken approaches may provide additional benefits. Included in these top hits are some synonyms, e.g. molnupiravir and eidd-2801 (a preclinical name for this drug), lopinavir-ritonavir as well as lopinavir/ritonavir and lpv/r. There are 11/14 instances of first mentions in the literature prior to 2020, all representing drugs that were repurposed for COVID-19. Thus, we note that with our methodology, it is also possible to automatically identify instances of drug repurposing from the literature.
Limitations
The embeddings-based approach to predicting research impact achieves AUC-Extreme10% values of 0.77 for the reference setup. While an AUC in the range between 0.7 and 0.8 is considered as acceptable [26], this does not mean that our method can, with acceptable reliability, assess the research impact of any article. Instead, our research was aimed at evaluating the ability of machine learning models to identify the most promising research.
In our work, we use the widely used Word2vec method to generate embeddings. The newer alternative approaches, such as BERT [27], have the disadvantage that they generate different word embeddings for each context. The use of contextual embeddings might improve predictive performance, but the adoption of stable word embeddings empowers our method to deliver simple explanations. In future work, we plan to extend the method by supporting contextual embeddings, such as those generated by BERT or Large Language Models, such as LLama [28].
In many applications, it is useful for features to correspond to multiword sequences such as bigrams or trigrams. Naturally, by defining $A:=(w,w^{\prime})$ for an entity-pair $w,w^{\prime}$, we can compute $\textit{SCORE}(A)$ using (3). Sorting the entities according to their scores can reveal which connections of words have a high potential to predict the high citability of a paper. Processing $n$-grams can have significance, especially in scientific areas such as medicine or material engineering that often use multi-word expressions to denote concepts.
In our work, we predict whether the paper will belong to the lowly/highly cited category, which may appear to be too coarse-grained compared to directly predicting the citation counts when it comes to identifying the ground-breaking research. However, even in this binary classification framework, the confidence of the prediction to the highly cited class correlates with actual research impact as shown through our experiments. Nevertheless, the methodology could be further extended by directly predicting citation counts. While in our previous work [5], we report on such an approach having lower predictive performance than building a binary classifier as done in the present article, the availability of specific counts of predicted citations could have various applications.
Conclusions
This article presents two interconnected contributions. Our original aim was to devise a supervised framework for predicting which research will become impactful. However, when we started working on this, we did not find a classification approach that would be both embedding-based and directly explainable. As the first contribution, we describe a new feature importance algorithm SMER that can explain a supervised logistic regression model trained on an embeddings-based representation by applying the model to documents corresponding to embedding word vectors for single words. We first theoretically proved that this explanation approach has 100% fidelity with respect to the underlying classification model. Next, we experimentally demonstrated the superiority of embeddings-based models over the bag-of-words representation in terms of predictive performance and the quality of explanations if word importance is calculated with SMER. In another experiment, we showed that SMER produces a better global feature importance ranking than would follow from averaged LIME scores.
These experiments have shown the merits of the method quantitatively, but does it also generate useful predictions? Our second contribution is a new methodology for identifying specific newly published but potentially highly impactful entities as well as complete research articles. We provide a detailed analysis of the significance of the top-predicted drugs, vaccines and viruses. The predictions are also validated through bibliometric analysis, which shows that the method correctly identified research written in some of the best journals by leading scientists. Note that while most scientometric research relies on bibliometric factors such as journal quality, number of authors, etc., our training methodology was based solely on the content of the articles as reflected in the article abstracts, and the bibliometric information is used purely for validation purposes. Therefore our method is less susceptible to introducing institutional, gender, and other biases into the machine learning models.
Declarations
Availability of code and data
Code is available from the authors at https://github.com/beranoval/PRECOG/tree/main. The code for Experiment 3 is at https://github.com/adrianakubecova/LIME_DP/blob/main/Lime.ipynb. The supplementary analysis comparing SMER with LIME on the IMDB dataset is at https://github.com/adrianakubecova/LIME_DP/blob/main/Lime_imdb.ipynb.
Competing interests
The authors have no potential conflicts of interest to disclose.
Funding
The work of M. Černý was supported by Czech Science Foundation under project 22-19353S. The work of L. Beranová Dvořáčková was supported by grant IGA 16/2022 “PRECOG: Predicting REsearch COncepts of siGnificance”. T. Kliegr was supported by the long-term support for research activities of FIS VSE. The authors would like to Karel Douda for acquiring the citation counts.
Author contributions
L.D. data processing and analysis, V.S. journal quality data and analysis, A.K. comparison between LIME and SMER and AOPC analysis, M.P.J. analysis targets and interpretation of results, T.K. and L.D. writing, M.Č. formalization, A.D. L.D. and T.K. methodology, T.K. conceptualization, approval of the final manuscript, all authors.

### v2

@@ Results
Of note is the fourth table (Table 1(d)) containing the highest predicted words by similarity to “Remdesivir”. This choice of the seed word followed a previous analysis [24] that detected topics related to COVID-19 and identified the antiviral drug Remdesivir as the dominant term for the drug topic. The word “Remdesivir” on its own has with $\textit{SCORE}(w)=0.961$, as shown in Table 1(d), also a high value of the predicted significance, although not as high as for imdevimab ($\textit{SCORE}(w)=0.997$), which was thus chosen as the representative for the drug topic.
Experiment 5: Qualitative evaluation of top-predicted papers
Unlike Experiment 4, which focuses on predictions of entities, this one focuses on the prediction of impactful papers. The predictions are validated through the bibliometric profile of the publishing journals and authors.
Datasets
We analyzed “young” papers published from the year 2021 which have the greatest potential (Score) according to the model.This experiment is based on the same data and predictive model as the previous experiment.
Evaluation strategy
The predictions of impactful papers are evaluated through bibliometric indicators.
Bibliometric indicators are widely accepted to correlate with research quality. For example, combined impact factor and citation counts were previously used to predict research impact [25]. As a supplementary approach for validating predictions, we use bibliometric information related to the journal where the paper appeared, and a bibliometric profile of its authors. We emphasize that this information has not been used in our training setup and is used only for validation.
The quality of the journal is measured through the ISI impact factor, which is computed as the citation counts to papers the journal published in 2019-2020 from papers published in the year 2021 in other journals with an impact factor divided by the number of papers published by the journal in the years 2019-2020. As a reference for a high-impact prediction, we use the median impact factor among the top 10% journals in the category “Health sciences” category (OECD category FORD 3030, data for the year 2021). To represent the research reputation of the authors, we use the maximum of the first and last authors’ Hirsch-index (H-index) values. The H-index is defined as the number of papers published by the author that have at least as many citations as this number. As a reference for a high-impact prediction, we use the median H-index of full professors, which was determined by analyzing the medical education journal editorial boards [26]. Previous research has found that the H-index was a good indicator of citation in the first few years after publication [25].
Results
Figure 4 analyses a subset of papers with the highest predicted impact (Score), which were published in journals with an impact factor. Among the most recent papers (published in 2022), the paper that appeared in a journal assigned the highest impact factor value (39.2) is also predicted to have the highest impact according to our method (SMER score). Out of the eight papers with an impact factor, seven appeared in journals assigned an impact factor higher than the median of 4.74 in the source corpus. Note that the impact factor or other journal metadata were not part of the training corpus. Similarly, the authors of most top-ranked papers included an author with an H-index exceeding the median for full professors on journal editorial boards. The high impact factors of journals in which the predicted impactful papers were published, as well as the track record of their authors, confirm the ability of our model to identify future research impact. The Digital Object Identifiers (DOIs) of the specific top-predicted papers are available in the supplementary material.
Figure 4: This figure shows the top papers by predicted impact (Score) published in 2021. The reference impact factor (blue line) represents the median among the top 10% journals in the category “Health sciences” category. The reference H-index (red line) corresponds to the median H-index of full professors. The attention score and tweets were retrieved from Altmetrics.com.
Discussion
In this section, we present a more detailed examination and discussion of the predicted impactful entities from Experiment 4. This analysis and interpretation were conducted with significant input and validation from our domain expert in biosciences (M.P.J.) to ensure the biological relevance.
An important property of embeddings is that they can provide predictions even for words that were not contained in the labelled training data for the classification model. This is not possible with models based on the BoW representation. An example is “nirmatrevil”, which has a high predicted score (0.99) according to our proposed embedding-based logistic regression model. Since this word only appeared for the first time in a paper published in the year 2022, it was not included in the abstracts of any of the papers on which the logistic regression model was trained.
In Table 1(c), we show results for the most important drug-related entities which are similar to the COVID-19 drug imdevimab. To summarize these findings, we note that of the top hits to this drug, seven of these top words are other COVID-19 antibody drugs with the antibody-drug designated suffix -mab. Potentially other methods designed for learning word subtokens could identify further impactful patterns based on word prefix and suffix patterns. However, we observe that the identified important features similar to imdevimab are perfectly categorized as antibody drugs, combinations, or more abstract related terms (’monoclonal’), and include many synonyms. Thus included are: regen-cov (representing the combination of casirivimab (#1) and imdevimab, ly-cov555 (synonym for bamlanivimab (#3)), regn-cov2 (synonym for casirivimab-imdevimab), ronapreve (synonym for casirivimab-imdevimab), ly-cov016 (also known as etesevimab (#2)), and regn10987 (synonym for imdevimab). In total, including identified synonyms, 8 out of the top important and similar words to imdevimab are synonyms of this drug or its combinations (e.g. the combinations casirivimab-imdevimab, regn-cov2 and ronapreve). The remaining words among the top hits are other COVID-19 antibody drugs, with monoclonal representing the general concept of monoclonal antibodies. With the exception of this general term, all other important words first appeared in the literature in 2020 or later, further validating their relevance for similarity to imdevimab, which is a new drug developed for COVID-19.
In Table 1(d), we also considered similar and important features related to the drug remdevisir, which has been used as a diagnostic word to assess word similarities for COVID-19 literature. In this case, all of the top important features were small-molecule COVID-19 drugs, with one exception being the antibody drug sarilumab approved for rheumatoid arthritis and repurposed for COVID-19. We note again that the suffix -ir, indicating an antiviral compound, is found for 8 of the 14 most important words similar to remdesivir, suggesting that subtoken approaches may provide additional benefits. Included in these top hits are some synonyms, e.g. molnupiravir and eidd-2801 (a preclinical name for this drug), lopinavir-ritonavir as well as lopinavir/ritonavir and lpv/r. There are 11/14 instances of first mentions in the literature prior to 2020, all representing drugs that were repurposed for COVID-19. Thus, we note that with our methodology, it is also possible to automatically identify instances of drug repurposing from the literature.
Limitations
In many applications, it is useful for features to correspond to multiword sequences such as bigrams or trigrams. Naturally, by defining $A:=(w,w^{\prime})$ for an entity-pair $w,w^{\prime}$, we can compute $\textit{SCORE}(A)$ using (4). Sorting the entities according to their scores can reveal which connections of words have a high potential to predict the high citability of a paper. Processing $n$-grams can have significance, especially in scientific areas such as medicine or materials engineering that often use multi-word expressions to denote concepts.
The area of multiword sequences is linked to the use of Word2vec method to generate embeddings in our experiments. The newer alternative approaches, such as BERT [28], generate different word embeddings for each context. The use of contextual embeddings might improve predictive performance, but the adoption of stable word embeddings empowers our method to deliver simple explanations.
Future work aimed at possible adaptations of our method for the use with contextual embeddings is outlined in the next subsection. Another possible source of performance improvements is the use of non-linear classifiers instead of logistic regression used in our experiments. In the last part of the limitations section, we discuss the possible benefits and disadvantages of using non-linear classifiers.
Incorporating contextual word embeddings into SMER
We have so far assumed a context-free embedding, i.e. a fixed mapping $w\in\mathcal{W}\longmapsto\widetilde{w}\in\mathbb{R}^{d}$. In contrast, a context-sensitive (or contextual) embedding allows the vector representation of a word to vary depending on the context in which it occurs.
More formally, let $A_{1},\dots,A_{n}$ be the training abstracts, and for each dictionary word $w\in\mathcal{W}$, define the set
$F_{w}:=\left\{(i,j)\ \middle|\ \text{the }j\text{-th word of }A_{i}\text{ is }w\right\}.$
For each $(i,j)\in F_{w}$, we define a context-sensitive embedding $\widetilde{w}(i,j)\in\mathbb{R}^{d}$. That is, $\widetilde{w}(i,j)$ may differ depending on the document $A_{i}$ and position $j$, even when referring to the same word $w$. To emphasize this dependence on context, we write $\widetilde{w}(i,j)$.
The main question is: How can context-sensitive embeddings be incorporated into the SMER methodology? There are at least two straightforward approaches.
Approach A: occurrence-average embedding
We can simply construct a universal embedding $\widetilde{w}$ by averaging the context-sensitive embeddings over all occurrences of $w$:
$\widetilde{w}:=\frac{1}{|F_{w}|}\sum_{(i,j)\in F_{w}}\widetilde{w}(i,j),\quad w\in\mathcal{W},$
and use the resulting vectors as input to the SMER method in the same way as with context-free embeddings.
Approach B: contextualized scoring via singleton abstraction
For each training abstract $A_{i}=(w_{i1},\dots,w_{i,N_{i}})$, we use the context-sensitive embeddings of its words at their respective positions $(i,j)$, and compute the abstract embedding as
$\widetilde{A}_{i}:=\frac{1}{N_{i}}\sum_{j=1}^{N_{i}}\widetilde{w}_{ij}(i,j),\quad i=1,\dots,n.$
These document embeddings $\widetilde{A}_{1},\dots,\widetilde{A}_{n}$ are then used to train the logistic regression model (2) to estimate $\widehat{\beta}_{0},\widehat{\beta}$. The resulting coefficients are global—they do not depend on any particular document or position.
To assign a score to a specific word in a fixed context, we construct an artificial one-word abstract $A_{n+1}:=(w)$, compute the corresponding context-sensitive embedding $\widetilde{w}(n{+}1,1)$, and define its score in this minimal context as
$\textit{SCORE}(w):=L(\widehat{\beta}_{0}+\widehat{\beta}^{T}\widetilde{w}(n{+}1,1)).$
This yields a score for each dictionary word in a consistent, minimal context, enabling a universal ranking across $\mathcal{W}$.
Theoretical correspondence and limitations
Approach A is, in essence, a natural extension of the original SMER method: it produces a single fixed embedding for each word by averaging its contextualized vectors, and then applies the standard SMER pipeline without further modification. As such, all theoretical properties established for SMER—most notably Proposition 1 (linearity of logits) and its consequences—remain valid under Approach A.
In contrast, Approach B operates with context-sensitive embeddings throughout, including during scoring. While it preserves the overall logic of SMER (computing document-level scores via logistic regression on average embeddings), the change in representation introduces subtle differences. Since the embeddings used at training and at scoring time may stem from different contexts, the theoretical properties of SMER are no longer automatically guaranteed. In particular, Propositions 1–3 have not been formally established under this setup and should therefore be interpreted with caution.
Extending SMER to nonlinear classifiers
The embedding-based approach to predicting research impact achieves AUC-Extreme10% values of 0.77 for the reference setup. While an AUC in the range between 0.7 and 0.8 is considered as acceptable [29], this does not mean that our method can, with acceptable reliability, assess the research impact of any paper. Instead, our research was aimed at evaluating the ability of machine learning models to identify the most promising research. Enhanced performance could be obtained by the use of non-linear classifiers. However, this may result in the loss of the viable interpretability properties as discussed in the following.
The SMER methodology fundamentally relies on the linearity of the decision function in the embedding space. Specifically, it assumes that the prediction for a document is based on a linear combination of its word embeddings—an assumption that enables key analytical properties such as the decomposition of document-level logits into word-level contributions (Proposition 1).
In practice, however, nonlinear classifiers such as neural networks or decision trees are commonly used to model more complex relationships in the data. A natural question arises: can the SMER scoring approach be meaningfully extended to such models?
One possible direction is to retain the general idea of SMER: scoring individual words by presenting them in a fixed, minimal context (e.g., a singleton abstract), while using a nonlinear classifier instead of a linear one. This would yield a score $\textit{SCORE}(w)$ that reflects the model’s prediction for that minimal input, even though the model no longer admits a linear decomposition.
However, this generalization comes with limitations. Without a linear decision surface, the prediction for a document can no longer be expressed as an average over its parts. Proposition 1 no longer holds, and fidelity cannot be guaranteed.
Nonetheless, applying SMER-style ideas in the nonlinear setting may still be useful in exploratory or diagnostic contexts, particularly when interpretability is desired but strict decomposability is not required. Such extensions may enable broader use of the SMER framework in modern NLP pipelines, but they require careful validation.
Conclusions
This paper presents two interconnected contributions. Our original aim was to devise a supervised framework for predicting which research will become impactful. However, when we started working on this, we did not find a classification approach that would be both embedding-based and directly explainable. As the first contribution, we describe a new feature importance algorithm SMER that can explain a supervised logistic regression model trained on an embedding-based representation by applying the model to documents corresponding to embedding word vectors for single words. We first theoretically proved that this explanation approach has 100% fidelity with respect to the underlying classification model. Next, we experimentally demonstrated the superiority of embedding-based models over the bag-of-words representation in terms of predictive performance and the quality of explanations if word importance is calculated with SMER. In another experiment, we showed that SMER produces a better global feature importance ranking than would follow from averaged LIME scores.
These experiments have shown the merits of the method quantitatively, but does it also generate useful predictions? Our second contribution is a new methodology for identifying specific newly published but potentially highly impactful entities as well as complete research papers. We provide a detailed analysis of the significance of the top-predicted drugs, vaccines and viruses. The predictions are also validated through bibliometric analysis, which shows that the method correctly identified research written in some of the best journals by leading scientists. Note that while most scientometric research relies on bibliometric factors such as journal quality, number of authors, etc., our training methodology was based solely on the content of the papers as reflected in the paper abstracts, and the bibliometric information is used purely for validation purposes. Therefore our method is less susceptible to introducing institutional, gender, and other biases into the machine learning models.
Declarations
Availability of code and data
Code is available from the authors at https://github.com/beranoval/PRECOG/tree/main.
Competing interests
The authors have no potential conflicts of interest to disclose.
Funding
The work of M. Černý was supported by Czech Science Foundation under project 23-06139S. The work of L. Beranová Dvořáčková was supported by grant IGA 16/2022 “PRECOG: Predicting REsearch COncepts of siGnificance”. T. Kliegr was supported by the long-term support for research activities of FIS VSE and the EU under agreement No 857446 (HeartBit 4.0). The authors would like to thank Karel Douda for acquiring the citation counts.
Author contributions
L.D. data processing and analysis, V.S. journal quality data and analysis, A.K. comparison between LIME and SMER and AOPC analysis, L.D. adaptation of AOPC analysis for SHAP and global surrogates, M.P.J. interpretation of results from the biosciences perspective, T.K. and L.D. writing, M.Č. formalization, A.K. L.D. and T.K. methodology, T.K. conceptualization, approval of the final manuscript, all authors.
Supplementary Material
An interesting observation is that Table 1(b) contains several terms related to the Moderna vaccine (“moderna”, “spikevax”, “modernas”, “mRNA-1273”). The two main terms from this list (by frequency) are “moderna” and “mRNA-1273”, both with a count of about 800 and very similar importance (SCORE) of 0.9548 and 0.9681, respectively. The high agreement in the score for these close synonyms, which, however, have distinct word vectors and were thus handled as two completely separate words, shows that the prediction results are consistent. This also indicates that when the count of papers where the term appears is low, the score estimates are less stable, as can be seen with “spikevax” (has a count of 41) and “modernas” (count of only 16), which both have lower scores.
@@ Conclusions
Table S1: Top words for selected topics. Sorted by the $\textit{SCORE}(w)$, Sim is the cosine similarity of the word vector to the word vector of the topic in the caption of each table, year indicates the earliest publication year in the corpus when the word appears.